\begin{thebibliography}{32}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Blocki et~al.(2012)Blocki, Blum, Datta, and Sheffet]{jlt2012blocki}
Jeremiah Blocki, Avrim Blum, Anupam Datta, and Or~Sheffet.
\newblock The johnson-lindenstrauss transform itself preserves differential privacy.
\newblock In \emph{Proceedings of the 2012 IEEE 53rd Annual Symposium on Foundations of Computer Science}, FOCS '12, pp.\  410–419, USA, 2012. IEEE Computer Society.
\newblock ISBN 9780769548746.
\newblock \doi{10.1109/FOCS.2012.67}.
\newblock URL \url{https://doi.org/10.1109/FOCS.2012.67}.

\bibitem[Brown et~al.(2024)Brown, Hayase, Hopkins, Kong, Liu, Oh, Perdomo, and Smith]{brown2024insufficientstatisticsperturbationstable}
Gavin Brown, Jonathan Hayase, Samuel Hopkins, Weihao Kong, Xiyang Liu, Sewoong Oh, Juan~C. Perdomo, and Adam Smith.
\newblock Insufficient statistics perturbation: Stable estimators for private least squares, 2024.
\newblock URL \url{https://arxiv.org/abs/2404.15409}.

\bibitem[Burkhardt et~al.(2025)Burkhardt, Keller, Orlandi, and Schwiegelshohn]{burkhardt2025distributeddifferentiallyprivatedata}
Jakob Burkhardt, Hannah Keller, Claudio Orlandi, and Chris Schwiegelshohn.
\newblock Distributed differentially private data analytics via secure sketching, 2025.
\newblock URL \url{https://arxiv.org/abs/2412.00497}.

\bibitem[Clarkson(2005)]{subgradient2005clarkson}
Kenneth~L. Clarkson.
\newblock Subgradient and sampling algorithms for l1 regression.
\newblock In \emph{Proceedings of the Sixteenth Annual ACM-SIAM Symposium on Discrete Algorithms}, SODA '05, pp.\  257–266, USA, 2005. Society for Industrial and Applied Mathematics.
\newblock ISBN 0898715857.

\bibitem[Clarkson \& Woodruff(2015)Clarkson and Woodruff]{mestimator2015clarkson}
Kenneth~L. Clarkson and David~P. Woodruff.
\newblock \emph{Sketching for <italic>M</italic>-Estimators: A Unified Approach to Robust Regression}, pp.\  921--939.
\newblock Society for Industrial and Applied Mathematics, 2015.
\newblock \doi{10.1137/1.9781611973730.63}.
\newblock URL \url{https://epubs.siam.org/doi/abs/10.1137/1.9781611973730.63}.

\bibitem[Clarkson \& Woodruff(2017)Clarkson and Woodruff]{input2012clarkson}
Kenneth~L. Clarkson and David~P. Woodruff.
\newblock Low-rank approximation and regression in input sparsity time.
\newblock \emph{J. ACM}, 63\penalty0 (6), January 2017.
\newblock ISSN 0004-5411.
\newblock \doi{10.1145/3019134}.
\newblock URL \url{https://doi.org/10.1145/3019134}.

\bibitem[Clarkson et~al.(2016)Clarkson, Drineas, Magdon-Ismail, Mahoney, Meng, and Woodruff]{fast2016clarkson}
Kenneth~L. Clarkson, Petros Drineas, Malik Magdon-Ismail, Michael~W. Mahoney, Xiangrui Meng, and David~P. Woodruff.
\newblock The fast cauchy transform and faster robust linear regression.
\newblock \emph{SIAM Journal on Computing}, 45\penalty0 (3):\penalty0 763--810, 2016.
\newblock \doi{10.1137/140963698}.
\newblock URL \url{https://doi.org/10.1137/140963698}.

\bibitem[Coleman \& Shrivastava(2021)Coleman and Shrivastava]{onepass2021coleman}
Benjamin Coleman and Anshumali Shrivastava.
\newblock A one-pass distributed and private sketch for kernel sums with applications to machine learning at scale.
\newblock In \emph{Proceedings of the 2021 ACM SIGSAC Conference on Computer and Communications Security}, CCS '21, pp.\  3252–3265, New York, NY, USA, 2021. Association for Computing Machinery.
\newblock ISBN 9781450384544.
\newblock \doi{10.1145/3460120.3485255}.
\newblock URL \url{https://doi.org/10.1145/3460120.3485255}.

\bibitem[Dwork et~al.(2006)Dwork, McSherry, Nissim, and Smith]{dwork2006privacy}
Cynthia Dwork, Frank McSherry, Kobbi Nissim, and Adam Smith.
\newblock Calibrating noise to sensitivity in private data analysis.
\newblock In \emph{Proceedings of the Third Conference on Theory of Cryptography}, TCC'06, pp.\  265–284, Berlin, Heidelberg, 2006. Springer-Verlag.
\newblock ISBN 3540327312.
\newblock \doi{10.1007/11681878_14}.
\newblock URL \url{https://doi.org/10.1007/11681878_14}.

\bibitem[Feldman et~al.(2009)Feldman, Fiat, Kaplan, and Nissim]{private2009feldman}
Dan Feldman, Amos Fiat, Haim Kaplan, and Kobbi Nissim.
\newblock Private coresets.
\newblock In \emph{Proceedings of the Forty-First Annual ACM Symposium on Theory of Computing}, STOC '09, pp.\  361–370, New York, NY, USA, 2009. Association for Computing Machinery.
\newblock ISBN 9781605585062.
\newblock \doi{10.1145/1536414.1536465}.
\newblock URL \url{https://doi.org/10.1145/1536414.1536465}.

\bibitem[Feldman et~al.(2017)Feldman, Xiang, Zhu, and Rus]{private2017feldman}
Dan Feldman, Chongyuan Xiang, Ruihao Zhu, and Daniela Rus.
\newblock Coresets for differentially private k-means clustering and applications to privacy in mobile sensor networks.
\newblock In \emph{Proceedings of the 16th ACM/IEEE International Conference on Information Processing in Sensor Networks}, IPSN '17, pp.\  3–15, New York, NY, USA, 2017. Association for Computing Machinery.
\newblock ISBN 9781450348904.
\newblock \doi{10.1145/3055031.3055090}.
\newblock URL \url{https://doi.org/10.1145/3055031.3055090}.

\bibitem[Gordon et~al.(2006)Gordon, Zhu, and Wang]{gordon2006lad}
Michael~D. Gordon, Ji~Zhu, and Li~Wang.
\newblock { Regularized Least Absolute Deviations Regression and an Efficient Algorithm for Parameter Tuning }.
\newblock In \emph{Sixth International Conference on Data Mining (ICDM'06)}, pp.\  690--700, Los Alamitos, CA, USA, December 2006. IEEE Computer Society.
\newblock \doi{10.1109/ICDM.2006.134}.
\newblock URL \url{https://doi.ieeecomputersociety.org/10.1109/ICDM.2006.134}.

\bibitem[Johnson \& Lindenstrauss(1984)Johnson and Lindenstrauss]{jlt1984jlt}
William Johnson and Joram Lindenstrauss.
\newblock Extensions of lipschitz maps into a hilbert space.
\newblock \emph{Contemporary Mathematics}, 26:\penalty0 189--206, 01 1984.
\newblock \doi{10.1090/conm/026/737400}.

\bibitem[Ledoux \& Talagrand(1991)Ledoux and Talagrand]{Ledoux1991ProbabilityIB}
Michel Ledoux and Michel Talagrand.
\newblock Probability in banach spaces: Isoperimetry and processes.
\newblock In \emph{Probability in Banach Spaces: Isoperimetry and Processes}, 1991.
\newblock URL \url{https://api.semanticscholar.org/CorpusID:118526268}.

\bibitem[Li et~al.(2019)Li, Liu, Sekar, and Smith]{free2019tian}
Tian Li, Zaoxing Liu, Vyas Sekar, and Virginia Smith.
\newblock Privacy for free: Communication-efficient learning with differential privacy using sketches.
\newblock \emph{CoRR}, abs/1911.00972, 2019.
\newblock URL \url{http://arxiv.org/abs/1911.00972}.

\bibitem[Liu et~al.(2024)Liu, Mao, Zhang, and Zhang]{liu2024frappe}
Weidong Liu, Xiaojun Mao, Xiaofei Zhang, and Xin Zhang.
\newblock Efficient sparse least absolute deviation regression with differential privacy.
\newblock \emph{Trans. Info. For. Sec.}, 19:\penalty0 2328–2339, January 2024.
\newblock ISSN 1556-6013.
\newblock \doi{10.1109/TIFS.2023.3349054}.
\newblock URL \url{https://doi.org/10.1109/TIFS.2023.3349054}.

\bibitem[Liu et~al.(2021)Liu, Kong, and Oh]{liu2021differentialprivacyrobuststatistics}
Xiyang Liu, Weihao Kong, and Sewoong Oh.
\newblock Differential privacy and robust statistics in high dimensions, 2021.
\newblock URL \url{https://arxiv.org/abs/2111.06578}.

\bibitem[Meng \& Mahoney(2013)Meng and Mahoney]{low2013meng}
Xiangrui Meng and Michael~W. Mahoney.
\newblock Low-distortion subspace embeddings in input-sparsity time and applications to robust linear regression.
\newblock In \emph{Proceedings of the Forty-Fifth Annual ACM Symposium on Theory of Computing}, STOC '13, pp.\  91–100, New York, NY, USA, 2013. Association for Computing Machinery.
\newblock ISBN 9781450320290.
\newblock \doi{10.1145/2488608.2488621}.
\newblock URL \url{https://doi.org/10.1145/2488608.2488621}.

\bibitem[Munteanu et~al.(2021)Munteanu, Omlor, and Woodruff]{pmlr-v139-munteanu21a}
Alexander Munteanu, Simon Omlor, and David Woodruff.
\newblock Oblivious sketching for logistic regression.
\newblock In Marina Meila and Tong Zhang (eds.), \emph{Proceedings of the 38th International Conference on Machine Learning}, volume 139 of \emph{Proceedings of Machine Learning Research}, pp.\  7861--7871. PMLR, 18--24 Jul 2021.
\newblock URL \url{https://proceedings.mlr.press/v139/munteanu21a.html}.

\bibitem[Munteanu et~al.(2023)Munteanu, Omlor, and Woodruff]{munteanu2023almost}
Alexander Munteanu, Simon Omlor, and David Woodruff.
\newblock Almost linear constant-factor sketching for \${\textbackslash}ell\_1\$ and logistic regression.
\newblock In \emph{The Eleventh International Conference on Learning Representations}, 2023.
\newblock URL \url{https://openreview.net/forum?id=gu-SC0dpkvw}.

\bibitem[Nelson \& Nguyên(2013)Nelson and Nguyên]{osnap2013nelson}
Jelani Nelson and Huy~L. Nguyên.
\newblock Osnap: Faster numerical linear algebra algorithms via sparser subspace embeddings.
\newblock In \emph{2013 IEEE 54th Annual Symposium on Foundations of Computer Science}, pp.\  117--126, 2013.
\newblock \doi{10.1109/FOCS.2013.21}.

\bibitem[Sarlos(2006)]{rp2006sarlos}
Tamas Sarlos.
\newblock Improved approximation algorithms for large matrices via random projections.
\newblock In \emph{2006 47th Annual IEEE Symposium on Foundations of Computer Science (FOCS'06)}, pp.\  143--152, 2006.
\newblock \doi{10.1109/FOCS.2006.37}.

\bibitem[Sheffet(2017)]{dpols2017sheffet}
Or~Sheffet.
\newblock Differentially private ordinary least squares.
\newblock In \emph{Proceedings of the 34th International Conference on Machine Learning - Volume 70}, ICML'17, pp.\  3105–3114. JMLR.org, 2017.

\bibitem[Sheffet(2019)]{pmlr-v98-sheffet19a}
Or~Sheffet.
\newblock Old techniques in differentially private linear regression.
\newblock In Aurélien Garivier and Satyen Kale (eds.), \emph{Proceedings of the 30th International Conference on Algorithmic Learning Theory}, volume~98 of \emph{Proceedings of Machine Learning Research}, pp.\  789--827. PMLR, 22--24 Mar 2019.
\newblock URL \url{https://proceedings.mlr.press/v98/sheffet19a.html}.

\bibitem[Sohler \& Woodruff(2011)Sohler and Woodruff]{subspace2011sohler}
Christian Sohler and David~P. Woodruff.
\newblock Subspace embeddings for the l1-norm with applications.
\newblock In \emph{Proceedings of the Forty-Third Annual ACM Symposium on Theory of Computing}, STOC '11, pp.\  755–764, New York, NY, USA, 2011. Association for Computing Machinery.
\newblock ISBN 9781450306911.
\newblock \doi{10.1145/1993636.1993736}.
\newblock URL \url{https://doi.org/10.1145/1993636.1993736}.

\bibitem[Upadhyay(2017)]{upadhyay2017differentiallyprivatelinearalgebra}
Jalaj Upadhyay.
\newblock Differentially private linear algebra in the streaming model, 2017.
\newblock URL \url{https://arxiv.org/abs/1409.5414}.

\bibitem[Varshney et~al.(2022)Varshney, Thakurta, and Jain]{varshney2022sgd}
Prateek Varshney, Abhradeep Thakurta, and Prateek Jain.
\newblock (nearly) optimal private linear regression for sub-gaussian data via adaptive clipping.
\newblock In Po-Ling Loh and Maxim Raginsky (eds.), \emph{Proceedings of Thirty Fifth Conference on Learning Theory}, volume 178 of \emph{Proceedings of Machine Learning Research}, pp.\  1126--1166. PMLR, 02--05 Jul 2022.
\newblock URL \url{https://proceedings.mlr.press/v178/varshney22a.html}.

\bibitem[Wang \& Xu(2022)Wang and Xu]{wang2022differentiallyprivateell1normlinear}
Di~Wang and Jinhui Xu.
\newblock Differentially private $\ell_1$-norm linear regression with heavy-tailed data, 2022.
\newblock URL \url{https://arxiv.org/abs/2201.03204}.

\bibitem[Wang(2018)]{wang2018revisitingdifferentiallyprivatelinear}
Yu-Xiang Wang.
\newblock Revisiting differentially private linear regression: optimal and adaptive prediction \& estimation in unbounded domain, 2018.
\newblock URL \url{https://arxiv.org/abs/1803.02596}.

\bibitem[Woodruff(2014)]{sketch2014woodruff}
David~P. Woodruff.
\newblock Sketching as a tool for numerical linear algebra.
\newblock \emph{Found. Trends Theor. Comput. Sci.}, 10\penalty0 (1–2):\penalty0 1–157, October 2014.
\newblock ISSN 1551-305X.
\newblock \doi{10.1561/0400000060}.
\newblock URL \url{https://doi.org/10.1561/0400000060}.

\bibitem[Woodruff \& Zhang(2014)Woodruff and Zhang]{woodruff2014subspaceembeddingsellpregressionusing}
David~P. Woodruff and Qin Zhang.
\newblock Subspace embeddings and $\ell_p$-regression using exponential random variables, 2014.
\newblock URL \url{https://arxiv.org/abs/1305.5580}.

\bibitem[Zhao et~al.(2022)Zhao, Qiao, Redberg, Agrawal, El~Abbadi, and Wang]{dpsketch2022zhao}
Fuheng Zhao, Dan Qiao, Rachel Redberg, Divyakant Agrawal, Amr El~Abbadi, and Yu-Xiang Wang.
\newblock Differentially private linear sketches: efficient implementations and applications.
\newblock In \emph{Proceedings of the 36th International Conference on Neural Information Processing Systems}, NIPS '22, Red Hook, NY, USA, 2022. Curran Associates Inc.
\newblock ISBN 9781713871088.

\end{thebibliography}

\documentclass[10pt]{article} 
\usepackage[preprint]{tmlr}

\usepackage{amsmath,amsfonts,bm}

\def\eqref#1{equation~\ref{#1}}

\def\1{\bm{1}}

\DeclareMathAlphabet{\mathsfit}{\encodingdefault}{\sfdefault}{m}{sl}
\SetMathAlphabet{\mathsfit}{bold}{\encodingdefault}{\sfdefault}{bx}{n}

\usepackage{hyperref}
\usepackage{url}
\usepackage{algorithm,algpseudocode}
\usepackage{amsthm}
\usepackage{bbm}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}{Lemma}

\newtheorem{definition}{Definition}

\def \notes{1}

\ifnum\notes=1

    \newcommand{\dnnote}[1]{{\footnote{{\color{magenta} {\bf DN:} #1}}}}
    
    \newcommand{\sdnote}[1]{{\footnote{{\color{blue} {\bf SD:} #1}}}}
        
    \newcommand{\adnote}[1]{{\footnote{{\color{magenta} {\bf AD:} #1}}}}

\else

    \newcommand{\dnnote}[1]{}
    \newcommand{\sdnote}[1]{}
    \newcommand{\adnote}[1]{}
    
\fi

\title{Private Sketches for Linear Regression}

\author{\name Shrutimoy Das \email shrutimoydas@iitgn.ac.in \\
      \addr Department of Computer Science and Engineering\\
      Indian Institute of Technology Gandhinagar
      \AND
      \name Debanuj Nayak \email dnayak@bu.edu \\
      \addr Boston University
      \AND
      \name Anirban Dasgupta \email anirbandg@iitgn.ac.in\\
      \addr Department of Computer Science and Engineering \\
      Indian Institute of Technology Gandhinagar
      }

\def\month{MM}  
\def\year{YYYY} 
\def\openreview{\url{https://openreview.net/forum?id=XXXX}} 

\begin{document}

\maketitle

\begin{abstract}
Linear regression is frequently applied in a variety of domains, some of which might contain sensitive information. This necessitates that the application of these methods does not reveal private information. Differentially private (DP) linear regression methods, developed for this purpose, compute private estimates of the solution. These techniques typically involve computing a  noisy version of the solution vector. Instead, we propose releasing private sketches of the datasets, which can then be used to compute an approximate solution to the regression problem. This is motivated by the \emph{sketch-and-solve} paradigm, where the regression problem is solved on a smaller sketch of the dataset instead of on the original problem space. The solution obtained on the sketch can also be shown to have good approximation guarantees to the original problem. Various sketching methods have been developed for improving the computational efficiency of linear regression problems under this paradigm. We adopt this paradigm for the purpose of releasing private sketches of the data. We construct differentially private sketches for the problems of least squares regression, as well as  least absolute deviations regression. We show that the privacy constraints lead to sketched versions of regularized regression. We compute the bounds on the regularization parameter required for guaranteeing privacy. The availability of these private sketches facilitates the application of commonly available solvers for regression, without the risk of privacy leakage.
\end{abstract}

\section{Introduction}
Differentially private  (DP) linear regression methods aim to compute the solution to the problem privately. Given a data matrix $X \in \mathbb{R}^{n \times d}$ and a response vector $y \in \mathbb{R}^{n},$ the linear regression problem is to compute $\beta^* = \underset{\tilde{\beta}\in \mathbb{R}^d}{\text{argmin}}||X\tilde{\beta} - y||_p^q,$ where $p=q=1$ denotes $\ell_1$ (least absolute deviations) regression and $p=q=2$ denotes $\ell_2$ (least squares) regression. DP linear regression algorithms estimate the solution $\beta^*$ privately. Most of the literature on DP linear regression focusses on  $\ell_2$ regression, while \citet{wang2022differentiallyprivateell1normlinear,liu2024frappe} proposed algorithms for DP $\ell_1$ norm regression.

Since the solution for  $\ell_2$ regression has a closed form expression, $\beta^* = (X^{\top}X)^{-1}X^{\top}y,$  existing approaches include computing private estimates of $X^{\top}X$ and $X^{\top}y$ \citet{wang2018revisitingdifferentiallyprivatelinear},  using random projections for the privately estimating $\beta^*$ \citet{dpols2017sheffet}, estimating the solution using private gradient descent type algorithms \citet{varshney2022sgd}, as well as exponential mechanism type approaches \citet{liu2021differentialprivacyrobuststatistics}. \citet{brown2024insufficientstatisticsperturbationstable} proposes estimating the covariance matrix on a ``good'' set  of data (discarding points with high leverage scores as well as high residuals)  and perturbing the least squares estimate. For the case of $\ell_1$ regression,  \citet{liu2024frappe} proposes an iterative algorithm for estimating the solution which includes a kernel density estimation step for the purpose of introducing noise. \citet{wang2022differentiallyprivateell1normlinear} proposes using the exponential mechanism for computing the solution privately.

Each of these methods, except \citet{dpols2017sheffet}, involve multiple steps for estimating some property of the data matrix, which is then used for making the solutions private. In this paper, we propose releasing differentially private sketches of the data for solving the linear regression problem. We consider the classic \emph{sketch-and-solve} paradigm \citep{sketch2014woodruff} where an approximate solution to the regression problem is computed on a sketch of the data matrix. The sketched data matrix has much fewer rows than the original data matrix due to which computing the solution on the sketch is  computationally efficient. Various sketching algorithms exist that are able to given a good approximation for both $\ell_1$ and $\ell_2$ regression problems \cite{rp2006sarlos,osnap2013nelson,low2013meng,sketch2014woodruff,input2012clarkson,pmlr-v139-munteanu21a,munteanu2023almost}. While sketching has been utilized for tasks such as private low rank approximation  \citet{upadhyay2017differentiallyprivatelinearalgebra} and distributed private data analysis \citet{burkhardt2025distributeddifferentiallyprivatedata}, to the best of our knowledge, this is the first work that has looked at releasing private sketches for $\ell_1$ and $\ell_2$ regression. We note that \cite{dpols2017sheffet} utilizes random projections of the data matrix for computing the solution to $\ell_2$ regression privately.

An advantage of releasing private sketches of the data matrix is that the sketched regression problem can be solved using readily available solvers for both $\ell_1$ and $\ell_2$ regression, instead of having to estimate various properties of the data. Additionally, these private sketches can be queried infinitely, without any loss of privacy. The idea of releasing private sketches is closely related to the idea of releasing private coresets, which has been done for clustering \citep{private2009feldman,private2017feldman}. In this paper, we have not looked at private coresets for linear regression since coresets for regression generally depend on importance sampling from a data dependent distribution and are sensitive to changes in the dataset.

One of the most widely used sketching algorithms for $\ell_2$ regression is the Johnson Lindenstrauss transform (JLT) \citep{jlt1984jlt,rp2006sarlos}. In \citet{dpols2017sheffet}, it was shown that JLT itself preserves differential privacy when the data matrix satisfies certain conditions. 
\citet{pmlr-v98-sheffet19a} also noted that adding noise to the data matrix for the purpose of private linear regression results in a ridge regression problem where the regularization coefficient can be set such that the solution is differentially private. Note that this is unlike the standard regularized regression problems, where the regularization coefficient is set in order to the minimize the risk.  In this work, we  show that when noise is added in order to make the sketches private, solving the regression problem on the released private sketch is same as solving a sketched regularized regression problem (for both $p=1$ and $p=2$). We derive bounds for the regularization coefficient in such cases. However, when using the JLT for releasing private sketches, we show that the noise addition method can be altered such that the regression problem on the sketch is actually a  sketched version of an ``unregularized'' regression problem. 
When the sketching matrix is \texttt{CountSketch} or the sketching algorithm of \citet{munteanu2023almost}, we still have to deal with a regularized regression problem. 

\paragraph{Organization of the paper} In Section \ref{sec:prelims}, we introduce some definitions as well as the notations to be used in the paper. Section \ref{sec:l2_reg} proposes private sketches for the $\ell_2$ regression problem while Section \ref{sec:l1_reg} proposes sketches for the $\ell_1$ regression problem.

\section{Preliminaries}\label{sec:prelims}
We consider datasets of the form $(X,y)$ where $X \in \mathbb{R}^{n \times d}$ is the data matrix and $y \in \mathbb{R}^n$ is the response vector. In most of the discussions, we take $A = [X;y],$ where $y$ is appended to the columns of $X$ resulting in the $n \times (d+1)$ matrix $A.$ Two datasets $A, A'$ are \emph{neighbouring} if they differ in a single row. We assume that the $\ell_2$ norm of the rows of $A$ are bounded by $B.$ Also, the matrix $A$ is assumed to have full column rank.

Let $A  = U \Sigma V^{\top}$ be the singular value decomposition (SVD) of $A.$ We use the notation $\sigma_{min}(A)$ to denote the minimum singular value of $A.$ The $\ell_p$ regression loss is denoted by $||A\beta_{-1}||_p^q = ||X\beta - y||_p^q,$ where $\beta_{-1} = \begin{bmatrix}
    \beta \\
    -1
\end{bmatrix},\beta \in \mathbb{R}^d.$ The symbols $\mu, \nu$ denotes approximation error and failure probability respectively, such that $0 < \mu,\nu < 1.$ We denote the indicator function by $\mathbbm{1}{\cdot}$ The absolute value of a quantity is denoted by $|.|.$ The multivariate isotropic normal distribution with mean $0$ and variance $\sigma^2$ will be denoted by $\mathcal{N}(0, \sigma^2 I_d)$ where $I_d$ is the $d \times d$ identity matrix.

We define some of the concepts that will be used in the paper.

\begin{definition}[Differential Privacy \citet{dwork2006privacy}]
    A randomized algorithm $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private $((\epsilon,\delta)-DP)$ if for all neighbouring datasets $A,A'$ and for all possible outputs $\mathcal{E}$ in the range of $\mathcal{M},$
    \begin{align}\label{eqn:DP_def}
        Pr(\mathcal{M}(A') \in \mathcal{E}) \leq \exp(\epsilon)\cdot Pr(\mathcal{M}(A) \in \mathcal{E}) + \delta.
    \end{align}
\end{definition}

 \begin{definition}[$\ell_2$ sensitivity] The $\ell_2$ sensitivity of a function $f : \mathbb{R}^{n \times d}\rightarrow \mathbb{R}^{r \times d}$ is defined as
 $$\Delta(f) = \underset{\text{neighbours } A,A'}{\text{sup}}||f(A) - f(A')||_2.$$
 \end{definition}

 \begin{definition}[Gaussian Mechanism \citet{dwork2006privacy}] The Gaussian mechanism $\mathcal{M}$ with noise level $\sigma = \frac{\Delta(f)}{\epsilon}\sqrt{2 \ln(1.25/\delta)}$ is defined as
 $$\mathcal{M}(A) = f(A) + \mathcal{N}(0,\sigma^2 I_{d}).$$
 \end{definition}

The Gaussian mechanism is $(\epsilon, \delta)$- differentially private.

\begin{definition}[The Johnson-Lindenstrauss Lemma \citet{jlt1984jlt,rp2006sarlos}] Let $0 < \mu,\nu < 1$ and $S = \frac{1}{\sqrt{r}}R \in \mathbb{R}^{r \times n}$ such that $R_{ij} \sim \mathcal{N}(0,1)$ are independent random variables. If $r = \Omega(\frac{\log d/\nu }{\mu^2}),$ then for a $d$-element subset $V \subset \mathbb{R}^n,$ with probability at least $1 - \nu,$ we have
\begin{align}\label{eqn:JL_guarantee}
    (1- \mu) ||v||_2^2 \leq ||Sv||_2^2 \leq (1+ \mu) ||v||_2^2,
\end{align}
for all $v \in V.$ The matrix $S$ is the JL Transform.
\end{definition}


\begin{definition}[Subspace Embedding \citet{input2012clarkson}] A $\text{poly}(d,\mu^{-1}) \times n$ matrix $S$ is said to be a subspace embedding for $\ell_2$ for a fixed $n \times d$ matrix $X,$ if for all $v \in \mathbb{R}^d,$
$$(1 - \mu)||Xv||_2 \leq ||SXv||_2 \leq (1+ \mu)||Xv||_2,$$
with probability at least $9/10.$
\end{definition}

\citet{subspace2011sohler,woodruff2014subspaceembeddingsellpregressionusing} showed that subspace embeddings for the $\ell_1$ norm have $O(d \log d)$ dilation. The \emph{sketch and solve} paradigm  for linear regression \citet{sketch2014woodruff} consists of two steps :
\begin{enumerate}
    \item Compute a sketch , $SX,$ of the data matrix $X,$ where $S$ is the sketching matrix. The number of rows in $SX$ is less than in $X.$
    \item Compute $\beta' = \underset{\tilde{\beta}}{\text{argmin }}||SX\tilde{\beta} - Sy||_p^q$ on the smaller sketch.
\end{enumerate}

Previous works have shown that the solution obtained on the sketched data is a good  approximation to the original problem. For $\ell_2$ regression, it was shown in \citet{rp2006sarlos,input2012clarkson} that $||X\beta' - y||_2 \leq (1 + \mu)||X \beta^* - y||_2$ with high probability. Also, for the case of $\ell_1$ regression, \citet{munteanu2023almost} constructed a sketching matrix $S$ which was able to obtain an $O(1)$ approximation to the original problem.

\section{Private Sketching for \texorpdfstring{$\ell_2$}{l2} regression} \label{sec:l2_reg}
The Johnson Lindenstrauss transform (JLT) \citet{jlt1984jlt} has been shown to preserve $(\epsilon, \delta)$-differential privacy when used for publishing a sanitized covariance matrix \citet{jlt2012blocki} of the data matrix $A.$ A necessary condition for the privacy to be preserved is that  $\sigma_{min}(A)$  must be sufficiently large. Following this,  \citet{dpols2017sheffet,pmlr-v98-sheffet19a}  used the JLT for computing a private JL projection of $A.$ This private sketch was then used for computing the solution to the $\ell_2$ regression problem.

The method proposed by \citet{dpols2017sheffet} required that $\sigma_{min}(A) \geq w,$ for a threshold $w,$ such that the algorithm for computing the  sketches is differentially private. In case $\sigma_{min}(A) < w,$ a matrix $w . I_{d+1}$ is appended to the rows of
$A$ to get $\hat{A} = \begin{bmatrix}
    A \\
    w.I_{d+1}
\end{bmatrix} \in \mathbb{R}^{(n+d+1) \times (d+1)}$ such that the $\sigma_{min}(\hat{A}) \geq w.$ The effect of concatenating $w I_d$ is to shift the singular values by  $w$ such that $\sigma_{min}(\hat{A})$ becomes greater than or equal to $w$  \citep{jlt2012blocki}.

\begin{align} \label{eqn:w_inc_sigma}
\begin{split}
    \hat{A}^{\top}\hat{A} = A^{\top}A + w^2I_{d+1}  &= A^{\top}A + w^2 VV^{\top}\\
    & = V(\Sigma^2 + w^2 I_{d+1})V^{\top} \\
    & = V \sqrt{\Sigma^2 + w^2I_{d+1}} U^{\top}U \sqrt{\Sigma^2 + w^2I_{d+1}}V^{\top}.
\end{split}
\end{align}

When $\sigma_{min}(A) \geq w,$ let $S \in \mathbb{R}^{r \times n}$ be the JLT (for a suitably chosen $r$). The private sketch released in this case is $SA.$  Let  $\beta = \underset{\tilde{\beta} \in \mathbb{R}^{d}}{\text{argmin}}||SX\tilde{\beta} - Sy||_2^2$ Then, for $r = \Omega(\mu^{-2} d \log d),$  using the guarantees of JLT and  from Theorem $12$ of \citet{rp2006sarlos}, with probability at least $1/3,$ we have
\begin{align} \label{eqn:ra_jl_guarantee}
   ||SA\beta_{-1}||_2^2 \leq (1+\mu)||A\beta_{-1}||_2^2 = (1+\mu) ||X\beta - y||_2^2 \leq (1+\mu)^3 ||X \beta^* - y||_2^2.
\end{align}
Here, since no alterations have to be made in order to guarantee privacy, solving the regression problem on $SA$ is same as solving the problem on a sketch of the unregularized problem.

 In case $\sigma_{min}(A) < w,$ let $S = [S_1 ; S_2]$ such that $S_1 \in \mathbb{R}^{r \times n }$ and  $S_2 \in \mathbb{R}^{r \times (d+1)},$ for $r = \Omega(\mu^{-2} d \log d).$ The private sketch released is $
 S\hat{A}.$  The least squares problem now becomes $\underset{\tilde{\beta}_{-1}}{\text{min}}||S\hat{A}\tilde{\beta}_{-1}||_2^2.$ As noted in \citet{pmlr-v98-sheffet19a}, $||\hat{A}\tilde{\beta}_{-1}||_2^2 = ||A\tilde{\beta}_{-1}||_2^2 + w^2||\tilde{\beta}_{-1}||_2^2,$ i.e, it is an instance of the ridge regression problem where the parameter $w^2$ is set to preserve $(\epsilon,\delta)$-differential privacy.
 
 Then,  we get $||S\hat{A}\tilde{\beta}_{-1}||_2^2 \leq (1+\mu) ||\hat{A}\tilde{\beta}_{-1}||_2^2 = (1+\mu)\left(||A\tilde{\beta}_{-1}||_2^2 + w^2||\tilde{\beta}_{-1}||_2^2\right),$ by the JL guarantee. From \citet{rp2006sarlos} and the JL guarantee, for  $\beta = \underset{\tilde{\beta} \in \mathbb{R}^d }{\text{argmin}}||S_1X\tilde{\beta} - S_1y||_2^2,$ we have
\begin{align} \label{eqn:ra_hat_jl_guarantee}
\begin{split}
    ||S\hat{A}\beta_{-1}||_2^2 &\leq (1+\mu) ||\hat{A}\beta_{-1}||_2^2 \\
    & \leq (1+\mu)^3 ||\hat{A}\beta^*_{-1}||_2^2 \\
     & \leq (1+\mu)^3 \left(||A\beta^*_{-1}||_2^2 + w^2||\beta^*_{-1}||_2 \right) \\
     & \leq (1+\mu)^3 \left( ||X\beta^* - y||_2^2 + w^2||\beta^*_{-1}||_2^2 \right)
\end{split}
\end{align}
 with constant probability. In this case, since we have to append $w.I_{d+1}$ in order to guarantee privacy, solving the regression problem on $S\hat{A}$ is same as solving the problem on a sketch of a ridge regression problem  where the regularization coefficient is $w^2 =  \frac{8B^2}{\epsilon}\left( \sqrt{2 r \ln(8/\delta)} + 2 \ln(8/\delta) \right)$ \citep{dpols2017sheffet}. 

 
 


 \subsection{Private JL Sketch with no regularization}
Here, we show that when $\sigma_{min}(A) = \gamma.w,$ for $0 < \gamma <1$ and $\gamma$ is close to $1,$ then the sketching algorithm of \citet{dpols2017sheffet} can be modified such that solving the regression problem on the private sketch does not require solving a regularized least squares problem. 
We want to  form a matrix $\hat{A}$  such that  $\sigma_{min}(\hat{A}) \geq w.$ Let $Q \in \mathbb{R}^{(d+1) \times (d+1)}$ be a matrix such that, for some scalar $c,$ we append $cQ$ to get $\hat{A} = \begin{bmatrix}
    A \\
    cQ
\end{bmatrix}.$ Similar to Equation \ref{eqn:w_inc_sigma}, we want that concatenating $cQ$ increases the singular values of the resulting matrix $\hat{A}.$ Thus, we have $\hat{A}^{\top}\hat{A} = A^{\top}A + c^2 Q^{\top}Q.$ If $Q^{\top}Q = V\Sigma^2V^{\top},$ then we get $\hat{A}^{\top}\hat{A} = V \Sigma^2(1 + c^2)V^{\top}.$ This shifts the singular values of $A$ by $c^2 \Sigma^2.$ Thus, we set $Q = V \Sigma V^{\top}$ and $c = \sqrt{\frac{w^2}{\sigma_{min}(A)^2} - 1} = \sqrt{\frac{1}{\gamma^2}-1}.$ Also, note that $Q^{\top}Q = V \Sigma^2 V^{\top}  = A^{\top}A.$

Let $S = [S_1 ; S_2]$ where $S_1 \in \mathbb{R}^{r \times n}$ and $S_2 \in \mathbb{R}^{r \times (d+1)}$ for $r = \Omega(\mu^{-2} d \log d).$ The private sketch released in this case is $S\hat{A}.$ The least-squares problem on the released sketch then becomes $\underset{\tilde{\beta}_{-1}}{\text{min}}||S\hat{A}\tilde{\beta}_{-1}||_2^2.$ Let  $\beta = \underset{\tilde{\beta}}{\text{argmin}}||S_1X\tilde{\beta} - S_1y||_2^2.$ Then, using the guarantees of the JLT and Theorem $12$ of \citet{rp2006sarlos}, with constant probability, we have

\begin{align}\label{eqn:cq_guarantee}
\begin{split}
    ||S\hat{A}\beta_{-1}||_2^2 & \leq (1 + \mu)||\hat{A}\beta_{-1}||_2^2  \\
    & \leq (1 + \mu)^3 ||\hat{A}\beta^*_{-1}||_2^2 \\
    &= (1 + \mu)^3(||A\beta^*_{-1}||_2^2 + c^2||Q\beta^*_{-1}||_2^2) \\
    & = (1+\mu)^3(1+c^2)||A\beta^*_{-1}||_2^2 \\
    & = (1 + \mu)^3(1+c^2)||X \beta^* - y||_2^2
\end{split}
\end{align}
where the equality in the fourth step comes from  $||Q\beta^*_{-1}||_2^2 = \beta^{*\top}_{-1}Q^{\top}Q\beta^*_{-1} = \beta^{*\top}_{-1}A^{\top}A\beta^*_{-1} = ||A\beta^*_{-1}||_2^2.$
Thus, by concatenating $A$ with $cQ,$ the singular values are scaled up by a factor of  $1+c^2 = 1/\gamma^2.$ Solving the regression problem on the private sketch is same as solving an unregularized regression problem.

Algorithm \ref{alg:private_JL} presents the proposed method. The values of the input parameters are same as in \citet{dpols2017sheffet}.  The proof of privacy of Algorithm \ref{alg:private_JL} is same as in \citep{dpols2017sheffet}.

\begin{algorithm}
\caption{\label{alg:private_JL}Private JL Sketch}
\begin{algorithmic}[1]
    \Require A matrix $A \in \mathbb{R}^{n \times (d+1)}$ and a bound $B$ on the $\ell_2$-norm of any row of $A,$ \\
    privacy parameters: $\epsilon, \delta > 0$\\
    parameter $r$ denoting number of rows in the projected matrix
    \State Set $w$ s.t. $w^2 = \frac{8B^2}{\epsilon}\big( \sqrt{2 r \ln(8/\delta)} + 2 \ln(8/\delta) \big)$
    \State Sample $Z \sim Lap(4B^2/\epsilon)$ and let $\sigma_{min}(A)$ denote the smallest singular value of $A.$
    \If{$\sigma_{min}(A)^2 > w^2 + Z + \frac{4B^2 \ln(1/\delta)}{\epsilon}$}\label{if_cond}
        \State Sample a $(r \times n)$ matrix $S$ whose entries are i.i.d sampled from $\mathcal{N}(0,1)$
        \State \Return $SA$
    \Else
        \State $Q = V \Sigma V^{\top},$ where $A = U \Sigma V^{\top}$
        \State $c = \sqrt{\frac{w^2}{\sigma_{min}(A)^2} - 1}$
        \State Let $\hat{A} = \begin{bmatrix}
            A\\
            cQ
        \end{bmatrix}$ 
        \State Sample a $(r \times (n+d+1))$ matrix $S$ whose entries are sampled i.i.d from $\mathcal{N}(0,1)$
        \State \Return $S\hat{A}$
    \EndIf
\end{algorithmic}
\end{algorithm}

\subsection{Private CountSketch for \texorpdfstring{$\ell_2$}{l2} regression} \label{sec:l2_cs}
Here, we look at another type of sketching method, using the \texttt{CountSketch} matrix, and discuss the alterations such that the sketching algorithm becomes differentially private. The CountSketch matrix gives low-distortion subspace embeddings for the $\ell_2$ norm in a data oblivious manner, and has been shown to give $(1+\mu)$-approximate solution to the $\ell_2$ regression problem \citep{input2012clarkson,osnap2013nelson,low2013meng}. In \citet{dpsketch2022zhao}, it was shown that a noisy initialization (calibrated to the $\ell_2$ sensitivity) results in a differentially private count sketch for the purpose of frequency estimation. Releasing private sketches for various downstream  problems have also been proposed in \citep{free2019tian,onepass2021coleman}.

Let $S \in \{-1,0,1\}^{r \times n}$ be a count sketch matrix with $r \ll n.$ Each column of $S$ has a single non-zero entry the coordinate for which is chosen uniformly at random. Each non-zero entry is either $1$ with probability $1/2$ or $-1$ with probability $1/2.$ The $i$th row of the sketch, $(SA)_{i,:}$ sums up the rows of $A$ (multiplied by $\pm 1$) that are present at the indexes corresponding to the non-zeros of $S_{i,:},$ and every row of $A$ is chosen exactly once. Thus,  $(SA)_{i,:} = \underset{j : \mathbbm{1}\{S_{ij} \in \{-1,1\}\}}{\sum}(\pm1).A_{j,:}.$ Let $A'$ be  a neighbouring dataset that does not agree with $A$ on the $k$th row. Then, the $\ell_2$ sensitivity of computing $SA$ can be bounded as

\begin{align} \label{eqn:l2_sensitivity_CS}
    \Delta = ||SA - SA'||_2 = ||(SA)_k - (SA')_{k,:}||_2 = ||A_{k,:} - A'_{k,:}||_2 \leq 2B.
\end{align}

So, if we add a noise vector $\tilde{\eta} \sim \mathcal{N}(0, \frac{8B^2 \ln(1.25/\delta)}{\epsilon^2}I_{d+1})$ to each row of $SA,$ then the algorithm  releases $SA + \eta$ as the private sketch and it is $(\epsilon,\delta)$-differentially private owing to the Gaussian mechanism. Here, $\eta$ is a $r \times(d+1)$ matrix where each row is a noise vector $\tilde{\eta}_i, i = 1, \ldots, r.$ Then, the $\ell_2$ regression problem becomes $\underset{\beta_{-1}}{\text{min}}||(SA + \eta)\beta||_2^2.$  
But we cannot use Theorem $30$ of \citet{input2012clarkson} to give approximation guarantees for $||SA\beta_{-1}||_2^2.$ Hence, we introduce the Gaussian noise for preserving privacy differently.

Let $\eta$ be a $p \times (d+1)$ matrix such that each row of $\eta$ is sampled independently from $\mathcal{N}(0, \frac{8B^2 \ln(1.25/\delta)}{\epsilon^2}I_{d+1}).$ We concatenate $\eta$ to the rows of $A$ in order to get $\hat{A} = \begin{bmatrix}
    A \\
    \eta
\end{bmatrix}$ which is a $(n+p)\times(d+1)$ matrix. Let the $r \times (n+p)$ countsketch matrix be $S = [S_1 ; S_2]$ where $S_1 \in \{-1,0,1\}^{r \times n}$ and $S_2 \in \{-1,0,1\}^{r \times p}$ are also countsketch matrices. Then, the private sketch to be released is $S\hat{A}.$ The effect of $S_2 \eta$ is to add noise vectors to each of the $r$ rows of $S_1A.$ Now, we specify the number of rows $p$ of $\eta.$

In order to make the sketch $S\hat{A}$ private, we want to add the noise vectors from $\eta$ to each of the $r$ rows of $S_1A.$ Thus, we want that $S_2$ maps the rows of $\eta$ to each row of $S\hat{A}$ at least once. This is analogous to the Coupon Collector's problem in that we have to collect $r$ coupons at least once in order to win. The expected number of trials required is then $p = O(r \log r).$ Thus, the noise matrix $\eta$ is a $O(r \log r) \times (d+1)$ matrix.

Let $\beta = \underset{\tilde{\beta} \in \mathbb{R}^d}{\text{argmin}}||SX\tilde{\beta} - Sy||_2^2.$  Then, from the subspace embedding guarantees of the CountSketch matrix as well as $\ell_2$ regression approximation guarantees from \citet{low2013meng}, with constant probability we have

\begin{align} \label{eqn:l2_cs_error}
\begin{split}
||S\hat{A}\beta_{-1}||_2^2 & \leq (1+\mu)||\hat{A}\beta_{-1}||_2^2 \\
& \leq (1+\mu)^3 ||\hat{A}\beta^*_{-1}||_2^2 \\
& = (1 + \mu)^3 \left(||A\beta^*_{-1}||_2^2 + ||\eta \beta^*_{-1}||_2^2\right)\\
& = (1+ \mu)^3 \left( ||X \beta^* - y||_2^2 + ||\eta \beta^*_{-1}||_2^2 \right).
\end{split}
\end{align}
From the inequality at \ref{eqn:l2_cs_error}, we observer that solving the least squares regression problem on the private sketch $S\hat{A}$ results in solving a sketched ridge regression problem. In Theorem \ref{thm:l2_CS_guarantee}, we bound the regularization coefficient. 

\begin{theorem}\label{thm:l2_CS_guarantee}
    Let $S \in \{-1,0,1\}^{r \times (n + r \log r)}$ be the sketching matrix which has been described  above and let  $r = poly(d, \mu^{-1}).$ Also, let $\eta$ be a $r \log r\times(d+1)$ matrix where the $i$th row $\tilde{\eta}_i \sim \mathcal{N}(0, \frac{8B^2 \ln(1.25/\delta)}{\epsilon^2}I_{d+1}).$
    Then, solving the $\ell_2$ regression problem on $S\hat{A}$ is same as solving a sketched ridge regression problem where the regularization coefficient is 
    \begin{align} \label{l2_CS_add_bound}
        ||\eta\beta_{-1}||_2 \leq \frac{13B}{\epsilon}\sqrt{r \log r \ln(1.25/\delta)}||\beta_{-1}||_2
    \end{align}
    with constant probability.
\end{theorem}
\begin{proof}
    Let $\sigma = \frac{2B}{\epsilon} \sqrt{2 \ln(1.25/\delta)}.$ Each $\tilde{\eta}_i$ is sampled from $\mathcal{N}(0, \sigma^2I_{d+1}).$ So, we have $\tilde{\eta}_i^{\top}\beta_{-1} \sim \mathcal{N}(0, \sigma^2||\beta_{-1}||_2^2).$ We are interested in bounding $||\eta \beta_{-1}||_2 = \left( \underset{i=1}{\overset{r \log r}{\sum}}(\tilde{\eta}_i^{\top}\beta_{-1})^2\right)^{\frac{1}{2}},$ where $\eta\beta_{-1}$ is an $(r \log r)$-dimensional vector of i.i.d normal random variables. From \citet{Ledoux1991ProbabilityIB}, for $t \geq 0,$ we have 
    \begin{align} \label{eqn:gaussian_tail_bound}
        Pr(||\eta\beta_{-1}||_2 \geq t) \leq 4 \exp\left( -\frac{t^2}{8 \mathbbm{E}[||\eta \beta_{-1}||_2^2]}\right).
    \end{align}
    Here, $\mathbbm{E}[||\eta\beta_{-1}||_2^2] = \underset{i=1}{\overset{r\log r}{\sum}} \mathbbm{E}[(\tilde{\eta}_i^{\top}\beta_{-1})^2] = r \log (r) \sigma^2 ||\beta_{-1}||_2^2.$ Suppose $||\eta\beta_{-1}||_2 \geq t$ with probability at most $1/4.$ Then,
    \begin{align}\label{eqn:constant_prob_bound}
        \begin{split}
            &4 \exp \left( -\frac{t^2}{8 \mathbbm{E}[||\eta\beta_{-1}||_2^2]} \right)  \leq \frac{1}{4} \\
            \implies &\frac{t^2}{8 \mathbbm{E}[||\eta\beta_{-1}||_2^2]} \geq \ln(16) \\
            \implies &t  \geq \sqrt{8 \mathbbm{E}[||\eta\beta_{-1}||_2^2] \ln(16)} \approx \frac{13 B}{\epsilon}\sqrt{r \log(r) \ln(1.25/\delta)}||\beta_{-1}||_2.
        \end{split}
    \end{align}
\end{proof}

\section{Private Sketching for \texorpdfstring{$\ell_1$}{l1} regression}\label{sec:l1_reg}

An analogue of the JLT for the problem of $\ell_1$-norm is the Cauchy transform \citep{subspace2011sohler,fast2016clarkson}, though  it has weaker guarantees (lopsided subspace embedding) .  The Cauchy transform is used for speeding up the computation of a distribution that is used for sampling from the rows of $A.$ This algorithm was proposed in \citet{subgradient2005clarkson} to obtain a $(1 + \mu)$ approximation to the solution of the $\ell_1$ regression problem. However, this algorithm outputs a weighted subsample from the dataset itself,  which could lead to the leakage of private information.

Instead we look at a recently proposed oblivious linear sketching technique that  satisfies the conditions of a lopsided embedding and gives an $O(1)$ approximation to the solution of the $\ell_1$ regression problem \citep{munteanu2023almost}. Similar sketches have also been utilized in \citep{mestimator2015clarkson,pmlr-v139-munteanu21a}. The sketch is analogous to a concatenation of multiple \texttt{CountMinSketches.} Similar to Section \ref{sec:l2_cs}, we append a noise matrix to the rows of $A$ and release a sketch of this augmented matrix as the private sketch. We use the sketching algorithm of \citet{munteanu2023almost}.

\subsection{An Illustration}
For the purpose of illustration, we consider a simpler sketching matrix $S \in \{0,1\}^{r \times n}$ where each column of $S$ has a single entry set to $1.$ 
Also, for any vector $\beta_{-1},$ assume that $||SA\beta||_1 \leq  O(1) ||A\beta||_1,$ with constant probability. The $i$th row of $SA$ can be written as $(SA)_{i,:} = \sum_{j : \mathbbm{1}\{S_{ij}=1\}}A_{j,:}.$ Let $A'$ be the neighbouring dataset such that $A$ and $A'$ differ in the $k$th row. Then, the $\ell_2$ sensitivity of $SA$ can be bounded as
\begin{align}\label{eqn:l2_sensitivity}
 \Delta  =  ||SA - SA'||_2 = ||(SA)_{k,:} - (SA')_{k,:}||_2 = ||A_{k,:} - A'_{k,:}||_2 \leq ||A_{k,:}||_2 + ||A'_{k,:}||_2 \leq 2B.
\end{align}
Thus, if we add a noise vector $\tilde{\eta} \sim \mathcal{N}(0, \frac{8B^2 \ln (1.25/\delta)}{\epsilon^2} I_{d+1}),$ then the algorithm that samples $S$ and releases $SA,$ is $(\epsilon,\delta)$-DP, by the Gaussian mechanism. But, as discussed in Section \ref{sec:l2_cs}, we will not be able to give approximation guarantees on the sketched $\ell_1$ problem , $||(SA + \eta)\beta_{-1}||_1.$ Here, $\eta$ is a $r \times(d+1)$ noise matrix for which each row is $\tilde{\eta}_i, i = 1 , \ldots,r.$ 

Instead we append a $p \times (d+1)$ noise matrix $\eta$ to the rows of $A$ to get $\hat{A} = \begin{bmatrix}
    A \\
    \eta
\end{bmatrix}$ and release the sketched matrix $S\hat{A}$ as the private sketch. Using the Coupon Collector's argument as before, $\eta$ is a $O(r \log r) \times (d+1)$ matrix. Solving the $\ell_1$ regression problem on $S\hat{A}$ is same as solving a sketched regularized $\ell_1$ regression problem. Let $\beta = \underset{\tilde{\beta}}{\text{argmin}}||SX\tilde{\beta} - Sy||_1,$  where $S \in \{0,1\}^{r \times (n + r \log r)}$ \citep{gordon2006lad}. Then, with constant probability, we have

\begin{align} \label{eqn:dp_l1sketch_loss}
\begin{split}
    ||S\hat{A}\beta_{-1}||_1 &\leq O(1) ||\hat{A}\beta_{-1}||_1 \\
    & \leq O(1) ||\hat{A}\beta^*_{-1}||_1 \\
    &= O(1) \left( ||A\beta^*_{-1}||_1 + ||\eta\beta^*_{-1}||_1 \right) \\
    &= O(1) \left( ||X\beta^* -y||_1 + ||\eta\beta^*_{-1}||_1 \right).
\end{split}
\end{align}
 We will be using Lemma \ref{lem:l1_tail_bound} for deriving a bound on the regularization coefficient.

\begin{lemma} \label{lem:l1_tail_bound}
    Let $U = [u_1, \ldots, u_r]$ be a vector of independent and identically distributed (i.i.d) random variables with $u_i \sim \mathcal{N}(0,\sigma^2).$ Let $||U||_1 = \underset{i=1}{\overset{r}{\sum}}|u_i| = S_r.$ Then, with probability at least $\frac{3}{4},$ 
    \begin{align} \label{eqn:l1_tail_bound}
        ||U||_1 \leq r \sigma    
    \end{align}
\end{lemma}

\begin{proof}
    Since $u_i \sim \mathcal{N}(0,\sigma^2),$ the random variable $Q = |u_i|$ has a half normal distribution. For $t \in \mathbb{R},$ the moment generating function (MGF) of $Q$ is
    \begin{align} \label{eqn:mgf_Q}
        M_{Q}(t) = \mathbb{E}[\exp(t|u_i|)] = \exp(\frac{\sigma^2t^2}{2}) \text{erfc}(-\frac{\sigma t}{\sqrt{2}}) = \exp(\frac{\sigma^2t^2}{2})(1 + \text{erf}(\frac{\sigma t}{\sqrt{2}})),
    \end{align}
    using the property that $\text{erf}(-z) = -\text{erf}(z), \text{erf}(.)$ being the Gaussian error function. Since, $u_i$s are i.i.d random variables, so too are the $|u_i|$s. Then, for $t \in \mathbb{R}$ , the MGF of $S_r$ becomes
            $M_{S_r}(t) = \mathbb{E}[\exp(t S_r)] = \mathbb{E}[\exp(t \underset{i=1}{\overset{r}{\sum}}|u_i|)] = \underset{i=1}{\overset{r}{\prod}} \mathbb{E}[t \underset{i=1}{\overset{r}{\sum}}|u_i|)] = [M_{Q}(t)]^r.$
    Applying Chernoff bound on the sum of i.i.d random variables, we have
    \begin{align}\label{eqn:chernoff_bound}
    \begin{split}
        Pr(S_r \geq a) & \leq \underset{t  > 0}{\inf} \exp(-ta) \underset{i=1}{\overset{r}{\prod}} \mathbb{E}[\exp(t |u_i|)] \\
        & = \underset{t  > 0}{\inf} \exp(-ta)[\exp(\frac{\sigma^2 t^2}{2})(1 + \text{erf}(\frac{\sigma t}{\sqrt{2}}))]^r 
    \end{split}
    \end{align}

    Let $g(t) = \exp(-ta)[\exp(\frac{\sigma^2 t^2}{2})(1 + \text{erf}(\frac{\sigma t}{\sqrt{2}}))]^r.$ Then, we get
    \begin{align}\label{eqn:ln_g_t}
        \begin{split}
            \ln (g(t)) &= \ln (\exp(-ta)[\exp(\frac{\sigma^2 t^2}{2})(1 + \text{erf}(\frac{\sigma t}{\sqrt{2}}))]^r) \\
            &= -ta + \frac{r \sigma^2 t^2}{2} + r \ln (1 + \text{erf}(\frac{\sigma t}{\sqrt{2}})) \\
            & < -ta  + \frac{r \sigma^2 t^2}{2} + r \ln 2,  
        \end{split}
    \end{align}
    since $\text{erf}(.) \in (-1,1).$ The minimizer of $\ln(g(t))$ is at $t' = \frac{a}{r \sigma^2}.$ At the minimizer, we get $\ln(g(t')) = \ln (2^r) - \frac{a^2}{2r \sigma^2}$ which gives $g(t') = \exp(ln(2^r))\exp(-\frac{a^2}{2r\sigma^2}) = 2^r\exp(-\frac{a^2}{2r\sigma^2}).$ Thus,
    \begin{align}
        Pr(S_r \geq a) \leq 2^r\exp(-\frac{a^2}{2r\sigma^2}). 
    \end{align}
     If $Pr(S_r \geq a) \leq \frac{1}{4},$ then
     \begin{align}\label{eqn:l1_reg_bound}
         \begin{split}
             &2^r\exp(-\frac{a^2}{2r\sigma^2}) \leq \frac{1}{4} \\
             \implies & a \geq r \sigma .
         \end{split}
     \end{align}
\end{proof}
 This leads us to Lemma \ref{lem:add_err_bound_simple}.

 \begin{lemma}\label{lem:add_err_bound_simple}
     With constant probability,
     \begin{align} \label{eqn:simple_add_bound}
         ||\eta \beta_{-1}||_1 \leq \frac{2B r \log r \sqrt{2 \ln(1.25/\delta)}}{\epsilon} ||\beta_{-1}||_1.
     \end{align}
 \end{lemma}
\begin{proof}
    We have $||\eta \beta_{-1}||_1 = \underset{i=1}{\overset{r\log r}{\sum}}|\eta_i \beta_{-1}|,$ where $\eta_i$ is the $i$th row of $\eta.$ Also, each coordinate of $\eta_i$ is sampled i.i.d from  $\mathcal{N}(0, \frac{8B^2 \ln (1.25/\delta)}{\epsilon^2}.$ Thus, $\eta_i\beta_{-1} \sim \mathcal{N}(0, \frac{8B^2 \ln (1.25/\delta)}{\epsilon^2}||\beta_{-1}||_2^2).$ Therefore, $||\eta\beta_{-1}||_1$ is the sum of $r \log r$ i.i.d random variables each sampled from $\mathcal{N}(0,\frac{8B^2 \ln (1.25/\delta)}{\epsilon^2}||\beta_{-1}||_2^2).$ Hence, by Lemma \ref{lem:l1_tail_bound}, we have
    \begin{align} \label{eqn:add_bound_simple}
        ||\eta\beta_{-1}||_1 \leq \frac{2B r \log r \sqrt{2\ln (1.25/\delta)}}{\epsilon}||\beta_{-1}||_2 \leq \frac{2B r \log r \sqrt{\ln (1.25/\delta)}}{\epsilon}||\beta_{-1}||_1,
    \end{align}
    with probability at least $\frac{3}{4}.$
\end{proof}

Now, we consider the sketching algorithm of \citet{munteanu2023almost} and apply the modifications to make it private.

\subsection{Private \texorpdfstring{$\ell_1$ Sketch}{l1}}
The sketching algorithm given in \cite{munteanu2023almost} outputs a weak weighted sketch that obtains an $O(1)$ approximation to the $\ell_1$ regression problem. The sketching dimension  is $r = O(d^{1+c} \ln(n)^{3+5c}), 0< c \leq 1 $ for constant success probability. The sketching algorithm of \citet{munteanu2023almost} also outputs an oblivious weight vector $w.$


Since, we append a $r \log r \times (d+1)$ noise matrix to $A$ (to get the noise augmented matrix $\hat{A}$), the weight vector $w$ is also of length $r \log r.$ Note that the weight vector as well as the sketch is data oblivious.
We give a brief description of the sketching algorithm in \citet{munteanu2023almost}.  The sketching matrix $S$ consists of $O(\log n)$ levels of sketch as shown.

\begin{align} \label{eqn:l1_sketch}
    S = \begin{bmatrix}
        S_0 \\
        S_1 \\
        \vdots\\
        S_{h_m}
    \end{bmatrix}.
\end{align}
Here, $h_m = O(\log_{b}n)$ is the number of sketch levels, where $b \in \mathbb{R}$ is a branching parameter. The number of buckets(rows) at level $h$ is denoted by $N_h.$ The probability that any row $A_{i,:}$ is sampled at level $h$ is $p_h \propto 1/b^h.$ The rows of $A$ that are mapped to the same row of the sketch are added up. Also, the weight of any bucket at level $h$ is set to $1/p_h.$

The sampling probabilities decrease exponentially as the levels increase. At level $0,$ sketching with $S_0$ is a \texttt{CountMinSketch} of the entire dataset. It captures the elements that make a significant contribution to the objective, the so-called heavy hitters. For improving efficiency, the rows at level $0$ are mapped to $s \geq 1$ rows of the sketch. On the other hand, the sketch at level $h_m$ corresponds to a small uniform subsample of the data. For each sketch level $S_i,$ the probability of having a $1$ at any of the columns is $1/b^{i}.$  For a detailed analysis of the algorithm, readers may refer to \citep{munteanu2023almost}.

The private sketch, $S\hat{A},$ sums up the corresponding rows of $A$ and at least one of the noisy rows from $\eta.$ The details are given in Algorithm \ref{alg:dp_l1_sketch}. It is similar to the algorithm of \citet{munteanu2023almost} with a few alterations.

\begin{algorithm}
  \caption{\label{alg:dp_l1_sketch}DP oblivious sketching algorithm for $\ell_1$ regression.}
  \begin{algorithmic}[1]
    \Statex \textbf{Input:} Data $A \in \mathbb{R}^{n \times (d+1)}$, number of rows $r=N \cdot h_m + N_u$, parameters $b>1, s\geq 1$ where $N=s \cdot N'$ for some $N' \in \mathbb{N},$ privacy parameters $\epsilon,\delta,$ $\ell_2$ row norm bound $B$;
    \Statex \textbf{Output:} weighted Sketch $C=(SA, w) \in \mathbb{R}^{r \times d}$ with $r$ rows.;
    \State $\sigma = \frac{2Bh_m}{\epsilon}\sqrt{2 \ln(1.25/\delta)},$ where $h_m = O(\log_b n)$
    \State $\hat{A} = \begin{bmatrix}
        A \\
        \eta
    \end{bmatrix},$ where each row of $\eta$ is $\eta_i \sim \mathcal{N}(0, \sigma^2 I_{d+1}).$
    \For{$h=0\ldots h_m$} \Comment{construct levels $0, \dots h_m$ of the sketch}
 		\State initialize sketch $S\hat{A}_h=  0$ at level $h$ ;
 		\State initialize weights $w_h=b^h \cdot \mathbf{1} \in \mathbb{R}^{N}$ at level $h$;
 	\EndFor

 	\State set $w_0=\frac{w_0}{s}$; \Comment{adapt weights on level $0$ to sparsity $s$}
    
 	\For{$i=1\ldots n$} \Comment{sketch the data}
 
 	        \For{$l=1\ldots s$} \Comment{densify level $0$}
 	            \State draw a random number $B_i \in [N']$;
 		        \State add $\hat{A}_{i,:}$ to the $((l-1) \cdot N'+ B_i)$-th row of $SA_0$;
 		    \EndFor
 		    
 		    \State assign $\hat{A}_{i,:}$ to level $h\in[1, h_m-1]$ with probability $p_h=\frac{1}{b^h}$; 
 	        \State draw a random number $B_i \in [N]$; 
 		    \State add $\hat{A}_{i,:}$ to the $B_i$-th row of $S\hat{A}_h$;
 		    \State add $S\hat{A}_{i,:}$ to uniform sampling level $h_m$ with probability $p_{h_m}=\frac{1}{b^{h_m}}$; 
 		\EndFor
 	\State  Set $S\hat{A}=(S\hat{A}_0, S\hat{A}_1, \ldots S\hat{A}_{h_m} )$;
 	\State  Set $w= (w_0, w_1, \ldots w_{h_m} )$;
 	\State  \textbf{return} $C = (S\hat{A}, w)$;
  \end{algorithmic}
\end{algorithm}

In Algorithm \ref{alg:dp_l1_sketch}, a row can be sampled at most $h_m$ times. Therefore, the 
 $\ell_2$ sensitivity of the applying the sketch is at most $2 B \sqrt{h_m}.$ We can now state Theorem \ref{thm:dp_l1_sketch}.

 \begin{theorem}\label{thm:dp_l1_sketch}
     Algorithm \ref{alg:dp_l1_sketch} satisfies $(\epsilon,\delta)$-differential privacy and outputs a weighted sketch  $C = (S \hat{A},w).$ Solving the $\ell_1$ regression problem on $S\hat{A}$ is same as solving a sketched regularized $\ell_1$ regression problem where the regularization coefficient is 
     \begin{align}\label{eqn:main_dp_add_error}
         ||\eta \beta_{-1}||_1 \leq \frac{2B r \log r \sqrt{2  h_m \ln(1.25/\delta)}}{\epsilon} ||\beta_{-1}||_1.
     \end{align}
     with constant probability.
 \end{theorem}
 \begin{proof}
     The differential privacy guarantee comes from the application of the Gaussian mechanism. The $\ell_2$ sensitivity of the sketch is $2B\sqrt{h_m}.$ Applying Lemma \ref{lem:add_err_bound_simple}, we get the bound on the regularization coefficient.
 \end{proof}

\section{Conclusion and Future Work}
In this paper, we proposed a few techniques for releasing differentially private sketches of the dataset. We utilized the JL transform as well as sparse sketches for computing the sketches. We observed that in most of the cases, introducing noise in order to preserve privacy results in having to solve sketch regularized regression problems. The bounds on the regularization coefficients highlight the regularization effect for guaranteeing privacy. The coefficients could potentially be large, which could lead to a large drop in utility of the solutions on the private sketches. For the JL transform, we showed that it is possible to get unregularized regression formulations in certain cases. While for the $\ell_2$ regression, Algorithm \ref{alg:private_JL} works, we explored the feasibility of releasing private CountSketch as well.  A future direction of work is to come up with similar constructions for the other problems, especially $\ell_1$ regression, as well. Specifically, we would like to have a better noise addition mechanism such that the regularization coefficient does not become large.




\bibliography{main}
\bibliographystyle{tmlr}


\end{document}